hi, i need help how to solve the linear system by using eliminations....
8x-4y=-76
5x+2y=-16
people have tried to help me but i dont get it at all.
: hi, i need help how to solve the linear system by using eliminations....
8x-4y=-76
5x+2y=-16
people have tried to help me but i dont get it at all.

Answer by stanbon(19738)

solve the linear system by using eliminations....
8x-4y=-76
5x+2y=-16
-------------
Multiply thru the 2nd equation by 2:
8x-4y=-76
10x+4y=-32
-------------
Add the two equation to eliminate the y-terms.
18x = -108
x = -6
--------------
Substitute that into 5x+2y = -16 to solve for y:
5*-6 + 2y = -16
-30 + 2y = -16
2y = 14
y = 7
-----------------
Solution: x=-6; y=7
=======================
Cheers,
Stan H.
Question 167834: hi, i need help how to solve the linear system by using eliminations....
8x-4y=-76
5x+2y=-16
people have tried to help me but i dont get it at all.
: hi, i need help how to solve the linear system by using eliminations....
8x-4y=-76
5x+2y=-16
people have tried to help me but i dont get it at all.

Answer by salvyb(3)

8x-4y=-76_5x+2y=-16
Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 4.
8x-4y=-76_2*(5x+2y=-16)
Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 4.
8x-4y=-76_2*(5x+2y)=2(-16)
Multiply 2 by each term inside the parentheses.
_2*(5x+2y)=-32
Multiply 2 by each term inside the parentheses.
_(10x+4y)=-32
Remove the parentheses around the expression 10x+4y.
_10x+4y=-32
Add the two equations together to eliminate y from the system.
10x+4y=-32_ 8x-4y=-76_18x =-108
Divide each term in the equation by 18.
x=-6
Substitute the value found for x into the original equation to solve for y.
8(-6)-4y=-76
Multiply 8 by each term inside the parentheses.
-48-4y=-76
Move all terms not containing y to the right-hand side of the equation.
-4y=-28
Divide each term in the equation by -4.
y=7
This is the final solution to the independent system of equations.
x=-6
y=7